OpenStudy (anonymous):

A builder of tract homes reduced the price of a model by 12%. If the new price is $ 124,000 , what was its original price? How much can be saved by purchasing the model?

OpenStudy (anonymous):

Cute original price... let's call it x !!! So, the new price, 12400, this is the old price, discounted by 12%... 124000 = x - (12% of x) 124000 = x - 0.12x 124000 = x(1-0.12) 124000 = 0.88x Well, you get the idea... enjoy x :D

OpenStudy (anonymous):

Thank you, I kinda solved it a while ago when I was waiting for some to answer my question. Well we have the same solution. Thank you!!